### v1

# Five dimensional Weyl double copy

Abstract: In this paper, we generalize the Weyl double copy to five-dimensional spacetime. We show that a special class of five-dimensional type N vacuum solutions admits a special class of degenerate Maxwell field that squares to give the Weyl tensor. The five-dimensional Weyl double copy relation defines a scalar field that satisfies the source-free Klein-Gordon equation on the curved background.
1 Introduction
Duality is a very fascinating phenomenon in fundamental physics. In particular, the connection between gauge and gravity theories is at the center of attention of many researchers. The double copy relation is the most remarkable realization of the gauge/gravity duality, which is inspired by the Kawai-Lewellen-Tye relation [1] and formulated by the color/kinematics duality [2, 3] of the scattering amplitude at the quantum level. The double copy perspective offers a completely different way to interpret gravity from the gauge theory side and vice versa. The success of the double copy relation has also motivated investigations at the classical level connecting classical solutions of gauge and gravity theories which pioneers a classical double copy relation from the Kerr-Schild construction [4].
In this paper, we will concentrate on the second exact classical double copy, namely the Weyl double copy (WDC) [5], which connects the Weyl tensor of the gravity theory and the field strength tensor of the Maxwell theory. The gauge invariant nature of the WDC makes it very efficient to reveal the mathematical connections of the two theories and it has drawn intensive attentions from various perspectives [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. However, the investigations of the WDC relation are now restricted in four-dimensional spacetime. [^2] In three-dimensional spacetime, a new version of classical double copy has been proposed in [25, 26], which involves the Cotton tensor rather than the Weyl tensor. It is somehow mysterious to construct a WDC relation in higher dimensions considering its twistorial foundation [8, 10]. In contrast, the amplitude double copy relation does not rely on the dimension of the spacetime [2, 3] and the Kerr-Schild construction also holds for generic dimensions, see, e.g., concrete examples of the Kerr-Schild double copy in dimensions other than four [27, 28, 29, 30, 31, 32, 33]. Hence, there is definitely an urgent need for new perspectives on how to generalize the Weyl double copy to higher dimensions, which is the primary goal of this work.
In this paper, we initialize the study of a five-dimensional Weyl double copy relation. The construction of the four-dimensional WDC formula constricts that it is only valid for two classes of algebraically special spacetime, namely the type D and type N spacetime. We start from the simplest type in five dimensions, the type N vacuum solution of Einstein gravity in the Coley-Milson-Pravda-Pravdova (CMPP) classification [34, 35, 36, 37] in the null frame. The spinor formalism in five dimensions [38] naturally provides an algebraic construction of a possible five-dimensional WDC formula for the type N solution. However, it is very difficult to prove that the Weyl tensor of any type N solution admits a decomposition to the square of a degenerate Maxwell field (solution of the source-free Maxwell’s equation) and some scalar fields that satisfy the wave equation either on the curved background or on Minkowski spacetime. Nevertheless, we find a self-contained reduction of the type N Weyl tensor and the degenerate Maxwell field, where one can confirm that any solution of this special class admits a special class of degenerate Maxwell field that squares to give the Weyl tensor. Moreover, a complex scalar field is defined from this five-dimensional WDC relation and the scalar field satisfies the wave equation on the curved background. Hence, a concrete WDC relation is uncovered in five dimensions. We then present the WDC relation for a special class of the five-dimensional pp-wave and Kundt solutions.
2 Null frame and CMPP classification
In four-dimensional spacetime, the Newman-Penrose formalism [39] is very effective for addressing problems in General Relativity. In higher-dimensional spacetime, it is of convenience to construct a similar system, namely the null frame. The basis vectors are typically chosen as
$l=e_{0}=e^{1},\quad n=e_{1}=e^{0},\quad m_{i}=e_{i}=m^{i}=e^{i},\quad i=2,...,n-1,$ (1)
with two null vectors $l$ and $n$, and $n-2$ spacelike vectors $m_{i}$. They are required to satisfy the orthogonal conditions
$l\cdot l=n\cdot n=l\cdot m^{i}=n\cdot m^{i}=0,$ (2)
and the normalization conditions
$l\cdot n=1,\quad m^{i}\cdot m^{j}=\delta^{ij}.$ (3)
The spin coefficients are determined from the following defined quantities
$L_{\mu\nu}=\nabla_{\nu}l_{\mu},\quad N_{\mu\nu}=\nabla_{\nu}n_{\mu},\quad M_{\mu\nu}^{k}=\nabla_{\nu}m^{k}_{\mu}.$ (4)
Based on the above null frame, CMPP [34, 35, 36, 37] proposed an algebraic classification method for tensors. Similar to the four-dimensional case, the CMPP classification is based on the maximum boost order of tensors. For any tensor $T_{\mu\nu...\rho}$, it can be projected onto the chosen null frame, and the boost weight of the null frame component $T_{ab...c}\equiv e^{\mu}_{\,{a}}e^{\nu}_{\,{b}}...e^{\rho}_{\,{c}}T_{\mu\nu...\rho}$ is defined as the number of the index 00 minus the number of the index $1$ in the null frame component indices. The maximum boost weight of all non-zero null frame components is called as the boost order of the tensor $T$. It could be proved that the boost order of a tensor depends only on the choice of null direction $l$. The essence of the CMPP classification is to choose a null vector $l$ to minimizes the maximum boost weight and another null vector $n$ that maximizes the minimum boost weight.
In general, the Weyl tensor could be decomposed as
$\begin{split}&C_{\mu\nu\rho\sigma}=\ \underbrace{4C_{0i0j}n_{\{\mu}m_{\nu}^{i}n_{\rho}m_{\sigma\}}^{j}}_{\text{boost weight }+2}+\underbrace{8C_{010i}n_{\{\mu}l_{\nu}n_{\rho}m_{\sigma\}}^{i}+4C_{0ijk}n_{\{\mu}m_{\nu}^{i}m_{\rho}^{j}m_{\sigma\}}^{k}}_{+1}\\ &+\underbrace{4C_{0101}n_{\{\mu}l_{\nu}n_{\rho}l_{\sigma\}}+4C_{01ij}n_{\{\mu}l_{\nu}m_{\rho}^{i}m_{\sigma\}}^{j}+8C_{0i1j}n_{\{\mu}m_{\nu}^{i}l_{\rho}m_{\sigma\}}^{j}+C_{ijkl}m_{\{\mu}^{i}m_{\nu}^{j}m_{\rho}^{k}m_{\sigma\}}^{l}}_{0}\\ &\hskip 56.9055pt+\underbrace{8C_{101i}l_{\{\mu}n_{\nu}l_{\rho}m_{\sigma\}}^{i}+4C_{1ijk}l_{\{\mu}m_{\nu}^{i}m_{\rho}^{j}m_{\sigma\}}^{k}}_{-1}+\underbrace{4C_{1i1j}l_{\{\mu}m_{\nu}^{i}l_{\rho}m_{\sigma\}}^{j}}_{-2}.\end{split}$ (5)
where the notation $C_{\mu\nu\rho\sigma}=C_{\{\mu\nu\rho\sigma\}}\equiv\frac{1}{2}(C_{[\mu\nu][\rho\sigma]}+C_{[\rho\sigma][\mu\nu]})$ should be understood. Specifically, a type N spacetime in the CMPP classification is defined when its Weyl tensor have the following form
$C_{\mu\nu\rho\sigma}=4C_{1i1j}l_{\{\mu}m_{\nu}^{i}l_{\rho}m_{\sigma\}}^{j}\equiv 8\psi_{ij}l_{\{\mu}m_{\nu}^{i}l_{\rho}m_{\sigma\}}^{j}.$ (6)
In the null frame, the Maxwell tensor can be decomposed as
$F_{\mu\nu}=F_{0i}(n_{\mu}m^{i}_{\nu}-n_{\nu}m^{i}_{\mu})+F_{ij}m^{i}_{\mu}m^{j}_{\nu}+F_{01}(n_{\mu}l_{\nu}-l_{\mu}n_{\mu})+F_{1i}(l_{\mu}m^{i}_{\nu}-l_{\nu}m^{i}_{\mu}),$ (7)
where the boost weights of the right hand side terms are 1, 0, 0, -1, respectively. A degenerate (type N) Maxwell tensor in the CMPP classification should have the following form
$F_{\mu\nu}=F_{1i}(l_{\mu}m^{i}_{\nu}-l_{\nu}m^{i}_{\mu})\equiv\phi_{i}(l_{\mu}m^{i}_{\nu}-l_{\nu}m^{i}_{\mu}).$ (8)
3 5d WDC relation for type N solutions
The WDC relation is better appriciated in spinorial formalism [5, 7]. In four-dimensional spacetime, the WDC relation is interpreted by the decomposition of the Weyl spinor
$\psi_{ABCD}=\frac{3c}{S}\phi_{(AB}\phi_{CD)},$ (9)
where $\phi_{AB}$ is a totally symmetric 2-spinor defined from the Maxwell tensor $F_{\mu\nu}$ and $S$ is a scalar field that solves the wave equation. The WDC relation (9) can be naturally extended to higher dimensional theory with a spinorial description. Though, the spinorial formalism and the classification of spinors have been defined in 5d spacetime [38], the Bianchi identity of the Weyl tensor and the Maxwell equation are not known in spinorial formalism in 5d. So we need to convert the WDC relation in null frame to examine the 5d WDC relation.
A key relation in 4d to convert relations in spinorial formalism to tetrad formalism is the decomposition of the 2d Levi-Civita tensor
$\epsilon_{AB}=o_{A}\iota_{B}-o_{B}\iota_{A},\quad\epsilon_{A^{\prime}B^{\prime}}=\bar{o}_{A^{\prime}}\bar{\iota}_{B^{\prime}}-\bar{o}_{B^{\prime}}\bar{\iota}_{A^{\prime}}.$ (10)
In 5d spinorial formalism, a similar 2d tensor $\Omega_{AB}$ is defined to raise or lower the spinor index [38]. One can show that $\Omega_{AB}$ can be decomposed as
$\begin{split}&\Omega_{AB}=(k_{Aa}n_{Bb}-k_{Ba}n_{Bb})\epsilon^{ab}=k_{Aa}n_{B}^{\ a}-k_{Ba}n_{A}^{\ a},\\ &\Omega^{AB}=(k^{A}_{\ a}n^{B}_{\ b}-k^{B}_{\ a}n^{A}_{\ a})\epsilon^{ab}=k^{A}_{\ a}n^{Ba}-k^{B}_{\ a}n^{Aa},\end{split}$ (11)
where we would refer to [38] for the notations and definitions. One can verify that the decomposition in (11) satisfies all the properties that $\Omega_{AB}$ should obey. Then the Kronecker symbol can be written as
$\begin{split}\delta^{B}_{A}=\Omega_{AC}\Omega^{BC}=k_{Aa}n^{Ba}-n_{A}^{\ a}k^{B}_{\ a}.\end{split}$ (12)
For the 5d type N spacetime, the only non-zero Weyl scalar is $\psi_{ij}=C_{\mu\nu\rho\sigma}n^{\mu}m^{i\nu}n^{\rho}m^{j\sigma}$ in null frame [34, 35, 36, 37]. Equivalently, the only non-zero little group bi-spinors in the spinorial formalism for Weyl tensor of a type N solution is $\psi_{abcd}^{(4)}=C_{\mu\nu\rho\sigma}\sigma^{\mu\nu}_{\ \ AB}\sigma^{\rho\sigma}_{\ \ CD}n^{A}_{\ a}n^{B}_{\ b}n^{C}_{\ c}n^{D}_{\ d}$ [38], where
$\sigma^{\mu\nu}_{\ \ AB}=\frac{1}{2}(\gamma^{\mu}_{\ AC}\gamma^{\nu C}_{\ \ B}-\gamma^{\nu}_{\ AC}\gamma^{\mu C}_{\ \ B}),$ (13)
and the $\gamma^{\mu}_{\ AC}$ matrices are defined in [38].
A natural generalization of the WDC formula to 5d for the type N spacetime can be written as
$\psi_{ABCD}n^{A}_{a}n^{B}_{b}n^{C}_{c}n^{D}_{d}\,\propto\,\phi_{AB}\phi_{CD}n^{(A}_{a}n^{B}_{b}n^{C}_{c}n^{D)}_{d}.$ (14)
Here we do not specify the scalar field as the case in 4d. Because there are more non-zero components of the Weyl spinor in 5d even for the simplest case, the type N spacetime. In principle, there could be more scalar fields involved. Note that the 4d case of the WDC relation is only valid for type D and type N spacetime. There is only one (complex) non-zero component of the Weyl spinor for both cases. The 5d WDC formula for the Weyl scalar of the type N case is
$\psi_{abcd}^{(4)}\,\propto\,\phi_{(ab}^{(2)}\phi_{cd)}^{(2)},$ (15)
where $\phi_{ab}^{(2)}$ is the only non-zero little group bi-spinors for Maxwell tensor which is defined by $\phi_{ab}^{(2)}=F_{\mu\nu}\sigma^{\mu\nu}_{\ \ AB}n^{A}_{\ a}n^{B}_{\ b}$ [38]. More explicitly we have
$\begin{split}&\psi_{1111}^{(4)}=\frac{3c}{S_{1}}\phi_{11}^{(2)}\phi_{11}^{(2)},\qquad\psi_{2222}^{(4)}=\frac{3c}{S_{1}^{\prime}}\phi_{22}^{(2)}\phi_{22}^{(2)},\\ &\psi_{1112}^{(4)}=\frac{3c}{S_{2}}\phi_{11}^{(2)}\phi_{12}^{(2)},\qquad\psi_{2221}^{(4)}=\frac{3c}{S_{2}^{\prime}}\phi_{22}^{(2)}\phi_{21}^{(2)},\\ &\psi_{1212}^{(4)}=\frac{c}{S_{3}}[\phi_{11}^{(2)}\phi_{22}^{(2)}+2(\phi_{12}^{(2)})^{2}].\end{split}$ (16)
Here we have introduced two independent complex scalar fields $S_{1}$, $S_{2}$ ( $S_{1}^{\prime}$, $S_{2}^{\prime}$ are their complex conjugate) and one real scalar field $S_{3}$. To convert to the null frame, we need to apply the following relation
$\begin{split}\gamma^{\mu}_{\ AC}\gamma^{\nu C}_{\ \ B}n^{A}_{\ a}n^{B}_{\ b}&=-\gamma^{\mu}_{\ AD}n^{A}_{\ a}n^{Dd}\gamma^{\nu}_{EB}k^{E}_{\ d}n^{B}_{\ b}+\gamma^{\mu}_{\ AD}n^{A}_{\ a}k^{D}_{\ d}\gamma^{\nu}_{\ EB}n^{Ed}n^{B}_{\ b}\\ &=-\sqrt{2}n^{\mu}\delta_{a}^{\ d}\varepsilon^{\nu}_{\ db}-\sqrt{2}\varepsilon^{\mu}_{\ ad}n^{\nu}\delta^{D}_{\ b}\\ &=\sqrt{2}\varepsilon^{[\mu}_{\ ab}n^{\nu]},\end{split}$ (17)
where $\varepsilon^{\mu}_{\ ab}=k_{a}^{A}\gamma^{\mu}_{\ AC}n_{b}^{C}$ [38]. In particular,
$\varepsilon^{\mu}_{11}=\frac{1}{\sqrt{2}}(m^{2\mu}-im^{3\mu}),\quad\varepsilon^{\mu}_{22}=\frac{1}{\sqrt{2}}(m^{2\mu}+im^{3\mu}),\quad\varepsilon^{\mu}_{12}=\varepsilon^{\mu}_{21}=im^{4\mu},$ (18)
where $m^{2\mu},m^{3\mu},m^{4\mu}$ are three spacelike vectors in the null frame. Finally, we uncover the connections bewteen the little group bi-spinors and the spacetime tensors as
$\psi_{abcd}^{(4)}=C_{\mu\nu\rho\sigma}\sigma^{\mu\nu}_{\ \ AB}\sigma^{\rho\sigma}_{\ \ CD}n^{A}_{\ a}n^{B}_{\ b}n^{C}_{\ c}n^{D}_{\ d}=2C_{[\mu\nu][\rho\sigma]}\varepsilon^{\mu}_{\ ab}n^{\nu}\varepsilon^{\rho}_{\ cd}n^{\sigma},$ (19)
and
$\phi_{ab}^{(2)}=F_{\mu\nu}\varepsilon^{\mu}_{\ ab}n^{\nu}.$ (20)
Then the WDC formulas can be established in the null frame as
$\begin{split}&\psi_{22}-\psi_{33}-2i\psi_{23}=\frac{3c}{S_{1}}[(\phi_{2})^{2}-(\phi_{3})^{2}-2i\phi_{2}\phi_{3}],\\ &\psi_{22}-\psi_{33}+2i\psi_{23}=\frac{3c}{S_{1}^{\prime}}[(\phi_{2})^{2}-(\phi_{3})^{2}+2i\phi_{2}\phi_{3}],\\ &\psi_{34}+i\psi_{24}=\frac{3c}{S_{2}}(\phi_{3}\phi_{4}+i\phi_{2}\phi_{4}),\\ &\psi_{34}-i\psi_{24}=\frac{3c}{S_{2}^{\prime}}(\phi_{3}\phi_{4}-i\phi_{2}\phi_{4}),\\ &\psi_{44}=\frac{c}{S_{3}}[(\phi_{2})^{2}+(\phi_{3})^{2}+4(\phi_{4})^{2}].\end{split}$ (21)
The above equations present a natural extension of the WDC formulas to 5d for type N spacetime. However the formulas in (21) are simply algebraic constructions between a Weyl-like rank four tensor and an antisymmetric rank two tensor. The central dogma of the WDC relation for the present case is that if a type N vacuum solution always admits a degenerate Maxwell field that can square to give the Weyl tensor and the scalar field satisfies the wave equation. An alternative statement of the WDC relation is that if any degenerate Maxwell field on the curved vacuum type N background can always give the Weyl tensor and the scalar field satisfies the wave equation. We are not able to prove the central dogma for a generic type N spacetime from the Bianchi identity for the Weyl tensor and the Maxwell’s equation for the field strength. In the next section, we will adapt the WDC formulas (21) to realize a concrete Weyl double copy relation for a special class of type N solutions.
4 Reduced type N solution and the WDC relation
We now consider a 4d-like reduction of the 5d type N spacetime where we impose the following restrictions for the Weyl scalars
$\psi_{4i}=0.$ (22)
The traceless property of the Weyl tensor then yields $\psi_{22}=-\psi_{33}$. Correspondingly, the single and zeroth copies are required to be
$\phi_{4}=0,\quad S_{3}\rightarrow\infty,$ (23)
where the divergent scalar field $S_{3}$ is chosen to turn off the extra component from the doubling in 5d. We postpone commenting more on this curious point by the end of this section. A higher dimensional Goldberg-Sachs-like theorem [40, 41] guarantees that for the CMPP type N spacetime the Weyl aligened null direction is geodesic (but no longer shear-free in general). Thus we can affine parameterize the geodesic and then take an appropriate spin transformation and a null rotation to make other basis vectors parallely propagated along $l$. Consequently, the following spin coefficients will vanish in such arrangement
$L_{i0}=0,\quad L_{10}=0,\quad M^{i}_{j0}=0,\quad N_{i0}=0.$ (24)
The Bianchi identities in the null frame [36] are now reduced to
$\begin{split}&D\psi_{ij}=-\psi_{ik}L_{kj},\\ &\delta_{[k}\psi_{j]i}=-\psi_{il}M^{l}_{[jk]}+M^{l}_{i[j}\psi_{k]l}+\psi_{i[j}L_{k]1}+2L_{1[j}\psi_{k]i},\\ &L_{k[i}\psi_{j]k}=0,\quad L_{k[j}\psi_{m]i}+L_{i[m}\psi_{j]k}=0,\quad\psi_{i[k}\psi_{jm]}=0,\end{split}$ (25)
where $D$, $\Delta$, and $\delta_{i}$ are the directional derivatives associated to the basis vectors $l$, $n$, and $m_{i}$, respectively. The Bianchi identities in the reduced case yield that
$\begin{split}&L_{22}=L_{33},\quad L_{32}=-L_{23},\quad L_{i4}=L_{4i}=0,\\ &M^{2}_{44}=M^{3}_{44}=0,\quad M^{2}_{42}=M^{3}_{43},\quad M^{2}_{43}=-M^{3}_{42}.\end{split}$ (26)
With the above reduction, the Maxwell’s equation (the full equations in the null frame are listed in the Appendix) is now reduced to
$\begin{split}&D\phi_{i}=-\phi_{j}L_{ji},\\ &\delta_{j}\phi_{i}-\delta_{i}\phi_{j}=-\phi_{i}L_{1j}+\phi_{j}L_{1i}-\phi_{j}L_{i1}+\phi_{i}L_{j1}+\phi_{k}M^{k}_{ji}-\phi_{k}M^{k}_{ij},\\ &\delta_{i}\phi_{i}=-\phi_{i}L_{1i}+\phi_{i}L_{i1}\phi_{i}M^{i}_{jj}.\end{split}$ (27)
We introduce the following definitions
$\begin{split}&\Psi_{4}=-(\psi_{22}+i\psi_{23}),\quad\bar{\Psi}_{4}=-(\psi_{22}-i\psi_{23}),\\ &\varphi_{2}=-\frac{1}{\sqrt{2}}(\phi_{2}+i\phi_{3}),\quad\bar{\varphi}_{2}=-\frac{1}{\sqrt{2}}(\phi_{2}-i\phi_{3}),\\ &\delta=\frac{1}{\sqrt{2}}(\delta_{2}-i\delta_{3}),\quad\bar{\delta}=\frac{1}{\sqrt{2}}(\delta_{2}+i\delta_{3}),\quad\delta^{\prime}=i\delta_{4},\quad S=\frac{S_{1}}{c}.\end{split}$ (28)
Then, the Bianchi identity and Maxwell’s equation could be rewritten as
$\begin{split}&D\Psi_{4}=-(L_{22}+iL_{23})\Psi_{4},\\ &\delta\Psi_{4}=-\frac{1}{\sqrt{2}}[(2L_{12}+2M^{2}_{33}-L_{21})+i(L_{31}-2L_{13}+2M^{2}_{32})]\Psi_{4},\\ &\delta^{\prime}\Psi_{4}=[(M^{3}_{42}-2M^{3}_{24})+i(M^{2}_{42}+L_{41}-2L_{14})]\Psi_{4},\end{split}$ (29)
and
$\begin{split}&D\varphi_{2}=-(L_{22}+iL_{23})\varphi_{2},\\ &\delta\varphi_{2}=-\frac{1}{\sqrt{2}}[(L_{12}+M^{2}_{33}-L_{21})+i(L_{31}+M^{2}_{32}-L_{13})]\varphi_{2},\\ &\delta^{\prime}\varphi_{2}=[(M^{3}_{42}-M^{3}_{24})+i(L_{41}-L_{14}+M^{2}_{42})]\varphi_{2}.\end{split}$ (30)
The WDC formula in the reduced case is given in a more compact form by
$\Psi_{4}=\frac{1}{S}(\varphi_{2})^{2}.$ (31)
Now it is clear that the previously chosen divergent scalar field $S_{3}$ is to prevent the mixing term constructed from $\varphi_{2}\bar{\varphi}_{2}$ in the Weyl doubling. Such term is not involved at all in the 4d WDC relation. So it is reasonable to have a special treatment in the 5d WDC formula. Here we initialize the 5d investigation by simply adapting the WDC relation and imposing that there is no Weyl scalar associated to $\varphi_{2}\bar{\varphi}_{2}$ term.
Substituting the WDC relation (31) into the Bianchi’s identities and simplifying them with the Maxwell’s equation, we obtain the equations for the scalar field
$\begin{split}&D\log{S}=-(L_{22}+iL_{23}),\\ &\delta\log{S}=\frac{1}{\sqrt{2}}(L_{21}-iL_{31}),\\ &\delta^{\prime}\log{S}=M^{3}_{42}+i(M^{2}_{42}+L_{41}).\end{split}$ (32)
Following closely the treatment in [9], one can prove the following integrability conditions for the above differential equations,
$\begin{split}&(D\delta-\delta D)\log{S}=\frac{1}{\sqrt{2}}D(L_{21}-iL_{31})+\delta(L_{22}+iL_{23}),\\ &(D\delta^{\prime}-\delta^{\prime}D)\log{S}=\frac{1}{\sqrt{2}}D(L_{21}-iL_{31})+\delta(L_{22}+iL_{23}),\\ &(\delta\delta^{\prime}-\delta^{\prime}\delta)\log{S}=\delta(M^{3}_{42}+iM^{2}_{42}+iL_{41})-\frac{1}{\sqrt{2}}\delta^{\prime}(L_{21}-iL_{31}),\end{split}$ (33)
where the commutators of the operators $D$, $\delta$, $\delta^{\prime}$ in [42] and the Ricci identities in [43] are applied. Finally, it is straightforward to verify that the scalar field $S$ solves the Klein-Gordon equation in curved spacetime
$\Box S=(\nabla_{\mu}l^{\mu}\Delta+\nabla_{\mu}n^{\mu}D+\nabla_{\mu}m^{i\mu}\delta_{i}+D\Delta+\Delta D+\delta_{i}\delta_{i})S=0,$ (34)
where we have applied the equations in (32), the commutators of the operators and the Ricci identities. This completes the zeroth copy of the 5d WDC relation for the reduced type N solutions. In the next section, we will present two concrete examples of the 5d WDC relation.
5 Exact solutions in 5d
5.1 5d pp-wave solution
The line-element of the 5d PP-wave solution in the convention of the CMPP null frame is given by [44]
$\text{d}s^{2}=2\text{d}u[\text{d}v+H(u,x_{2},x_{3},x_{4})\text{d}u]+\text{d}x_{2}^{\ 2}+\text{d}x_{3}^{\ 2}+\text{d}x_{4}^{\ 2}.$ (35)
Adapted to the reduction of the previous section, one obtains
$H(u,x_{2},x_{3},x_{4})=f(u)x_{4}+G(u,x_{2},x_{3}),$ (36)
where $G(u,x_{2},x_{3})=g(u,z)+\bar{g}(u,\bar{z})$ and $z=x_{2}-ix_{3}$. We choose the same null frame system as [33]. Then, it is easy to show
$L_{22}=L_{23}=L_{21}=L_{31}=L_{41}=M^{2}_{42}=M^{3}_{42}=0.$ (37)
A generic solution to (32) for the 5d pp-wave solution is
$S=S(u,\bar{z}).$ (38)
The Weyl scalar is given by
$\Psi_{4}=\partial_{x_{2}}\partial_{\bar{z}}\bar{g}(u,\bar{z})=\partial_{\bar{z}}^{2}\bar{g}(u,\bar{z}),$ (39)
which allows us to determine the Maxwell scalar as
$\varphi_{2}=\sqrt{S(u,\bar{z})\partial_{\bar{z}}^{2}\bar{g}(u,\bar{z})}.$ (40)
We have checked directly that the Maxwell scalar satisfies the Maxwell’s equation. Hence we uncover a concrete WDC relation for the 5d pp-wave solution.
5.2 5d Kundt solution
The type N Kundt line-element is given by [45]
$\text{d}s^{2}=2\text{d}u[\text{d}v+H(u,v,x^{k})\text{d}u+W_{i}(u,v,x^{k})\text{d}x^{i}]+\delta_{ij}\text{d}x^{i}\text{d}x^{j},$ (41)
where $i,j,k=2,3,4$, and
$\begin{split}&H(u,v,x^{k})=\frac{v^{2}}{2y^{2}}+H^{(0)}(u,x^{i}),\quad y=\delta_{ij}x^{i}x^{j},\\ &W_{2}=-2\frac{v}{y},\quad W_{m}(u,v,x^{k})=x^{n}B_{nm}(u)+C_{m}(u),\quad m,n=3,4.\end{split}$ (42)
The reductions from $\psi_{24}=\psi_{34}=\psi_{44}=0$ yield that
$\begin{split}&H^{(0)}=f(u,x_{2},x_{3})+g(u,x_{2},x_{4}),\\ &g(u,x_{2},x_{4})=\frac{1}{2}B_{43}(u)^{2}x_{4}^{2}+g_{1}(u)x_{2}x_{4}+C_{3}(u)x_{4}+g_{2}(u,x_{2}).\end{split}$ (43)
Projecting the solution into the null frame system defined in [33], one can obtain
$L_{22}=L_{23}=L_{31}=L_{41}=M^{2}_{42}=M^{3}_{42}=0,\quad L_{21}=-\frac{1}{y}.$ (44)
A generic solution to (32) for the 5d Kundt solution is
$S=\frac{P(u,x_{2}+ix_{3})}{x_{2}}.$ (45)
Weyl scalar of the Kundt solution is given by
$\Psi_{4}=\frac{1}{(x_{2})^{3}}\bigg{[}(x_{2}+ix_{3})B_{43}(u)^{2}-iB_{43}(u)C_{4}(u)\\ -i\partial_{x_{3}}f(u,x_{2},x_{3})-x_{2}\partial_{x_{3}}^{2}f(u,x_{2},x_{3})+ix_{2}\partial_{x_{2}}\partial_{x_{3}}f(u,x_{2},x_{3})\bigg{]}.$ (46)
Then, one can recover the Maxwell scalar from taking the square-root $\varphi_{2}=\sqrt{S\Psi_{4}}$. It can be verified directly that the Maxwell scalar satisfies the Maxwell’s equation. Thus, a concrete WDC relation has been uncovered for the 5d Kundt solution.
6 Concluding remarks
In this paper, we establish a 5d WDC relating the curvature of a special class of the type N spacetime to an electromagnetic field strength. Our proposal, for the first time, extends the 4d WDC relation to higher dimensions. A higher-dimensional spinorial formalism [38] allows one to construct the possible WDC formula in a universal way in any dimension. Hence, our investigation can be in principle extended, with certain adaptions, to any dimension for a special class of type N or type D spacetime, which would significantly enlarge the scope of the WDC relation.
Acknowledgments
The authors thank Hong Lü for useful discussions and Ricardo Monteiro for useful correspondence on the integrability conditions. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grants No. 11935009, No. 12375066, and No. 12247103.
Appendix A Maxwell’s equation in the null frame
$DF_{01}+\delta_{i}F_{0i}+F_{0i}(M^{i}_{10}+M^{i}_{01}+M^{i}_{jj})-F_{01}(L_{10}+L_{jj})\\ -F_{0i}(M^{i}_{01}-N_{0i})-F_{ij}M^{j}_{0i}+F_{01}N_{00}-F_{1i}M^{i}_{00}=0,$ (47)
$\Delta F_{01}-\delta_{i}F_{1i}+F_{01}(N_{01}+N_{jj})-F_{1i}(M^{i}_{10}+M^{i}_{01}+M^{i}_{jj})\\ +F_{0i}M^{i}_{11}+F_{ij}M^{j}_{1i}+F_{01}L_{11}+F_{1i}(M^{i}_{10}-L_{1i})=0,$ (48)
$\Delta F_{0k}+\delta_{i}F_{ik}+DF_{1k}+F_{ok}(N_{01}+N_{jj})+F_{1k}(L_{10}+L_{jj})+F_{ik}(M^{i}_{10}+M^{i}_{01}+M^{i}_{jj})\\ +F_{0i}(M^{i}_{k1}-N_{ki})+F_{ij}M^{j}_{ki}+F_{01}(L_{k1}-N_{k0})+F_{1i}(M^{i}_{k0}-L_{ki})=0,$ (49)
$DF_{1i}+\Delta F_{i0}+\delta_{i}F_{01}=F_{01}L_{1i}+F_{j1}L_{ji}+F_{01}N_{0i}+F_{0j}N_{ji}+F_{1i}N_{00}+F_{ji}N_{j0}\\ +F_{10}M^{i}_{10}+F_{1j}M^{i}_{j0}+F_{i0}L_{11}+F_{ij}L_{j1}+F_{10}M^{i}_{01}+F_{j0}M^{i}_{j1},$ (50)
$DF_{ij}+\delta_{j}F_{0i}+\delta_{i}F_{j0}=F_{0i}L_{1j}+F_{ki}L_{kj}+F_{01}M^{i}_{0j}+F_{0k}M^{i}_{kj}+F_{1j}M^{i}_{00}+F_{0j}M^{i}_{10}\\ +F_{kj}M^{i}_{k0}+F_{i1}M^{j}_{00}+F_{i0}M^{j}_{10}+F_{ik}M^{j}_{k0}+F_{10}M^{j}_{0i}+F_{k0}M^{j}_{ki}+F_{j0}L_{1i}+F_{jk}L_{ki},$ (51)
$\Delta F_{ij}+\delta_{j}F_{1i}+\delta_{i}F_{j1}=F_{1i}N_{0j}+F_{ki}N_{kj}+F_{10}M^{i}_{1j}+F_{1k}M^{i}_{kj}+F_{0j}M^{i}_{11}+F_{1j}M^{i}_{01}\\ +F_{kj}M^{i}_{k1}+F_{i0}M^{j}_{11}+F_{i1}M^{j}_{01}+F_{ik}M^{j}_{k1}+F_{01}M^{j}_{1i}+F_{k1}M^{j}_{ki}+F_{j1}N_{0i}+F_{jk}N_{ki},$ (52)
$\delta_{k}F_{ij}=F_{i1}M^{j}_{0k}+F_{i0}M^{j}_{1k}+F_{il}M^{j}_{kl}+F_{1j}M^{i}_{0k}+F_{0j}M^{i}_{1k}+F_{lj}M^{i}_{lk},$ (53)
where the permutation of the indices $i,j,k$ for the last equation should be understood.